# Rise of the forsaken relics: connecting present-day stellar streams and phase-mixed galaxies to the Epoch of Reionization

### v1

@@ Appendix B Slope estimates from Bayesian modeling
In section 6.1, we use Bayesian models of Schechter functions to estimate the low-mass/faint-end slopes of the SMFs/UVLFs. This approach avoids binning, which is important because otherwise the slope estimates depend somewhat on the choice of binning scheme. The Bayesian approach described here produces more robust slope determinations.
@@
Our results are insensitive to the high-mass/bright-end limit of the integrations as long as our highest-mass/brightest progenitor galaxy lies within this limit. The low-mass/faint-end limit does affect our results significantly, as discussed in 6.1. We count the number of progenitors $n$ in a given sample to estimate the expected number $N$ using a Poisson distribution:
@@ Appendix C Bootstrapping Methodology
In section 6.1, we used bootstrapping to compute a central estimate of the low-mass/faint-end slope and the scatter across the 10 simulated halos. First, we use the mean and standard deviation of the posterior distribution of the $\alpha$ parameter to draw samples from a Gaussian distribution with this mean and standard deviation. The assumption of Gaussianity is supported by the posterior distribution of $\alpha$ shown in the left panel of Figure 16. Next, we use bootstrapping with replacement to select 10 halos at random such that we get 10 values of $\alpha$, and calculate their mean ($\mu_{\alpha}$) and standard deviation ($\sigma_{\alpha}$). Then we repeat these two steps 10000 times to get samples of $\mu_{\alpha}$ and $\sigma_{\alpha}$. Finally, we compute the median values of these samples, $\tilde{\mu}_{\alpha}$ and $\tilde{\sigma}_{\alpha}$, and define them as our central estimate of the low-mass/faint-end slope and the associated halo-to-halo scatter across the 10 halos. We report these values in Table 2. Figure 17 shows the distribution of our $\mu_{\alpha}$ and $\sigma_{\alpha}$ samples for the proto-MW/LG Schechter mass function, with the vertical blue line indicating the median values $\tilde{\mu}_{\alpha}$ and $\tilde{\sigma}_{\alpha}$, respectively.

### v2

@@ IX Summary
We would like to thank the reviewer for their insightful and valuable feedback which enhanced the clarity of the manuscript. AK would like to thank Zachary Langford for valuable discussions on Bayesian modeling. We would also like to thank Arpit Arora, Guochao Sun, and Nora Shipp for insightful suggestions, discussions, and comments that have improved the manuscript. We acknowledge use of the marmalade cluster at the University of Pennsylvania, on which the analysis of this work was done. AK and RES acknowledge support from NSF grants AST-2007232 and AST-2307787. RES further acknowledges support from NASA awards 22-ROMAN22-0055, 22-ROMAN22-0011, and 22-ROMAN22-0013, Simons Foundation grant 1018462, and Sloan Foundation award FG-2023-20669. AK and AL acknowledge support from the Charles E. Kaufman foundation through grant KA2022-129518. MBK acknowledges support from NSF grants AST-1910346, AST-2108962, and AST-2408247; NASA grant 80NSSC22K0827; HST-GO-16686, HST-AR-17028, HST-AR-17043, JWST-GO-03788, and JWST-AR-06278 from the Space Telescope Science Institute, which is operated by AURA, Inc., under NASA contract NAS5-26555; and from the Samuel T. and Fern Yanagisawa Regents Professorship in Astronomy at UT Austin.
@@ Appendix B Slope estimates from Bayesian modeling
In section VI.1, we use Bayesian models of Schechter functions to estimate the low-mass/faint-end slopes of the SMFs/UVLFs. This approach avoids binning, which is important because otherwise the slope estimates depend somewhat on the choice of binning scheme. The Bayesian approach described here produces more robust slope determinations.
@@
Our results are insensitive to the high-mass/bright-end limit of the integrations as long as our highest-mass/brightest progenitor galaxy lies within this limit. The low-mass/faint-end limit does affect our results significantly, as discussed in VI.1. We count the number of progenitors $n$ in a given sample to estimate the expected number $N$ using a Poisson distribution:
@@ Appendix C Bootstrapping Methodology
In section VI.1, we used bootstrapping to compute a central estimate of the low-mass/faint-end slope and the scatter across the 10 simulated halos. First, we use the mean and standard deviation of the posterior distribution of the $\alpha$ parameter to draw samples from a Gaussian distribution with this mean and standard deviation. The assumption of Gaussianity is supported by the posterior distribution of $\alpha$ shown in the left panel of Figure 16. Next, we use bootstrapping with replacement to select 10 halos at random such that we get 10 values of $\alpha$, and calculate their mean ($\mu_{\alpha}$) and standard deviation ($\sigma_{\alpha}$). Then we repeat these two steps 10000 times to get samples of $\mu_{\alpha}$ and $\sigma_{\alpha}$. Finally, we compute the median values of these samples, $\tilde{\mu}_{\alpha}$ and $\tilde{\sigma}_{\alpha}$, and define them as our central estimate of the low-mass/faint-end slope and the associated halo-to-halo scatter across the 10 halos. We report these values in Table 2. Figure 17 shows the distribution of our $\mu_{\alpha}$ and $\sigma_{\alpha}$ samples for the proto-MW/LG Schechter mass function, with the vertical blue line indicating the median values $\tilde{\mu}_{\alpha}$ and $\tilde{\sigma}_{\alpha}$, respectively.